# Evolving Assembly Code in an Adversarial Environment

### v1

@@
Abstract.
In this work, we evolve assembly code for the CodeGuru competition. The competition’s goal is to create a survivor—an assembly program that runs the longest in shared memory, by resisting attacks from adversary survivors and finding their weaknesses. For evolving top-notch solvers, we specify a Backus Normal Form (BNF) for the assembly language and synthesize the code from scratch using Genetic Programming (GP). We evaluate the survivors by running CodeGuru games against human-written winning survivors. Our evolved programs found weaknesses in the programs they were trained against and utilized them. In addition, we compare our approach with a Large-Language Model, demonstrating that the latter cannot generate a survivor that can win at any competition. This work has important applications for cyber-security, as we utilize evolution to detect weaknesses in survivors. The assembly BNF is domain-independent; thus, by modifying the fitness function, it can detect code weaknesses and help fix them. Finally, the CodeGuru competition offers a novel platform for analyzing GP and code evolution in adversarial environments. To support further research in this direction, we provide a thorough qualitative analysis of the evolved survivors and the weaknesses found.
@@ Abstract.
doi: 10.1145/nnnnnnn.nnnnnnnisbn: 978-x-xxxx-xxxx-x/YY/MMconference: The Genetic and Evolutionary Computation Conference 2024; July 14–18, 2024; Melbourne, Australiajournalyear: 2024ccs: Software and its engineering Assembly languagesccs: Software and its engineering Genetic programmingccs: Software and its engineering Source code generationccs: Software and its engineering Search-based software engineeringccs: Security and privacy Vulnerability managementccs: Computing methodologies Genetic programming
1. Introduction
@@
CodeGuru Xtreme (Leshem and Eyzenberg, 2012) is a coding competition where short 8086 assembly programs, called survivors, are loaded into a random address in a virtual computer memory arena. Their goal is to defeat all other survivors by staying the last program to run. An opponent is defeated when it runs an illegal command caused, e.g., by overwriting its memory. The screen-shot of the game screen is depicted in Figure 1. Each survivor gets a different color in the arena, representing the bytes it wrote to the shared memory. We elaborate on the game in Section 3.
@@
Figure 1. The CodeGuru Xtreme game. On the left are the survivors; on the center is the code of the selected survivor; and on the right is the arena, i.e., the memory status. Each survivor gets a different color in the arena, representing the bytes it wrote to the shared memory.
@@
In this work, we evolve winning survivors from scratch, i.e., from randomly generated assembly code, and without access to the source code of known survivors. For this task, we utilize Grammar-Guided Genetic Programming (G3P)—an evolutionary computation technique that incorporates GP principles, employs context-free grammar and operates directly with tree-based representations. G3P allows us to evolve assembly programs following grammar-type constraints. The goal of the individuals, embodied by the fitness function, is to overtake adversaries and win the game. The evolved code is represented using an Abstract Syntax Tree (AST) based on matching assembly BNF we defined, where the nodes and leaves are the BNF’s functions and terminals, respectively. Since our BNF and the operators we used are general and domain-agnostic, the approach applies to generating assembly programs for other domains and processors.
@@
No previous work has been done on the CodeGuru Xtreme game, except for an undergraduate project (Sysman and Leibu, 2009) and some early work on the “Core War” game (Andersen, 2001; Corno et al., 2003), which served as the basis for CodeGuru Xtreme (see Section 3). As elaborated in Section 2, some work has been done on the evolution of low-level languages (i.e., assembly and Java bytecode), and some work has utilized large-language models (LLM) to improve existing assembly code. Note that improving existing code is a simpler task than generating new code from scratch, as the former’s state space is much smaller (Petke et al., 2018; Banzhaf, 2018). To support that, we compared our approach to using LLM in Section 5.2, demonstrating that the latter fails to create a survivor that wins any competition, even when playing against simple survivors.
@@
The CodeGuru Xtreme competition has been running since 2005, with all past survivors publicly available. Thus, winning the competition is considerably tricky and requires, among other qualities, a good understanding of the 8086 assembly language.
@@
This work has implications for cyber-security. As demonstrated below, we utilize evolution to detect weaknesses in other survivors. In addition, understanding the assembly language is a necessity in some viruses. By modifying the fitness function, our approach can be used for detecting weaknesses in code and help in fixing them, detecting suspicious adversarial code, or, on the contrary, can be intended to avoid security mechanisms. Finally, our approach can evolve high-quality individuals with a small number of examples and without detailed information.
@@
The CodeGuru Xtreme competition provides a unique opportunity to analyze Genetic Programming (GP) and code evolution in adversarial environments. In order to encourage further research in this area, we conducted a comprehensive qualitative analysis of the evolved survivors and identified their strengths and weaknesses. This analysis sheds light on the effectiveness of the evolved code and provides valuable insights for future improvements and advancements in the field.
2. Related Work
@@
Several works have been done on low-level code evolution, some similar to assembly, like Verilog (a hardware description language) and Java bytecode. Karpuzcu (Karpuzcu, 2005) used Grammatical Evolution (GE) for evolving a simple program of a one-bit full adder. In spite of the strong bias they employed, the success achieved was only about 5.7%.
Orlov and Sipper (Orlov and Sipper, 2011) proposed a method for genetic improvement and repair of existing Java programs or any software that can be compiled into Java bytecode. Although Java bytecode resembles assembly, it has a simplified representation and does not have direct memory access. This contrasts assembly, which has very strong correspondence between its instructions, the architecture’s machine code instructions, and memory. Orlov and Sipper seeded the initial population with copies of a single hand-crafted individual and improved it over generations, while we evolve assembly code from scratch.
Rosin (Rosin, 2019) synthesized simple programs with loops from input/output examples. They targeted simplified low-level language similar to assembly, where each instruction consists of an opcode and a single operand.
@@
Some works focused on constrained assembly evolution—specific routines, predefined input and output tables, and manually written code parts. In (Serruto and Casas, 2017), multi-objective linear genetic programming is applied for the automatic generation of some specific assembly driver routines. The architecture used was an 8051 microcontroller assembly, which is similar to the 8086 we use. The evolved programs did not contain jump instructions because they could form infinite loops, in contrast to our wish to include them in the generated code. The results showed that the automatically generated microcontroller code for specific tasks can compete with a human programmer with a smaller code size or faster execution. We aim to recreate this result within an adversarial environment.
Similarly, (Ferrel and Alfaro, 2020) presented a methodology for writing Arduino board programs using an automatic generator of assembly language routines based on a cooperative co-evolutionary multi-objective linear genetic programming algorithm. They decomposed the problem into sub-components, generating about 73% of the program, and the remaining 27%, which are the main program and initial configuration routines, are manually written.In our case, we cannot break the goal of winning an opponent to sub-tasks without reading its code first, which we avoid. In addition, there is no clear way to represent a winning result in our case using an input-output table. We need to evolve from scratch, a winning program, including jumps and loops, completely automatically. The presented method (Ferrel and Alfaro, 2020) requires more general implementation and adaptations to match our use case, yet prove the ability to achieve it.
@@
Several articles compare recent program synthesis and evolution techniques. (Pantridge et al., 2017) compares Flash Fill (Microsoft Excel feature which uses version-space algebras to perform program synthesis from examples on string manipulation tasks), MagicHaskeller (synthesizes functional Haskell programs through an exhaustive search of programs with the correct type signatures. Uses Monte-Carlo algorithm to remove semantically equivalent programs from the search space), TerpreT (a probabilistic programming language that is designed for inductive program synthesis), and two forms of GP—PushGP (evolves programs in a Turing complete, stack-based language called Push) and Grammar Guided Genetic Programming (a GP that uses context-free grammar as a guideline for syntactically correct programs throughout the evolution). The comparison focused on the possibility of producing a program that passes all tests for all training and unseen testing inputs. It was done on problems of two types: the first is usually approached using machine code or assembly language (basic execution models problems), while the second is usually approached using high-level languages (general program synthesis benchmark suite). In the low-level field, the system that performed the best was TerpreT, while sadly, there were no gathered results for G3P. In the high-level field, both GP approaches outperformed the others.
A survey of recent developments in program synthesis with evolutionary algorithms (Dominik Sobania, 2021), found that the most influential approaches in the field are stack-based GP (usually PushGP), G3P, and Linear GP (uses registers for calculations and direct memory access). Encouraged by this review and by the fact that G3P has not been widely tested on low-level languages, we want to use G3P for the evolution of assembly programs.
Google DeepMind created AlphaDev, a deep-reinforcement learning agent trained to improve a sorting algorithm (Mankowitz et al., 2023). They trained AlphaDev with up to 16 Tensor Processing Units (TPU) v.3, with a total batch size of 1,024 per TPU core and for 1 million iterations. In practice, across all tasks, training took them, in the worst case, two days to converge. Aside from the high resource demands for representing and training the neural network, their approach assumes a given working program to improve. It is not designed to create programs from scratch.
A day after AlphaDev publication, a user on X (formerly Twitter) tried to use ChatGPT to optimize the same sorting algorithm (Papailiopoulos, 2023). As noted before, improving an existing code is simpler than generating one from scratch. Nevertheless, we evaluate this approach in Section 5.2 and test whether it is applicable to our domain.
@@
3. CodeGuru
@@
CodeGuru Xtreme (Leshem and Eyzenberg, 2012) is a coding competition based on “Core War”—a 1984 programming game created by D. G. Jones and A. K. Dewdney (Dewdney, 1984). In CodeGuru Xtreme, short 8086 assembly programs (at most 512 bytes long) of 16-bit commands, called survivors, are loaded into random space on a virtual computer memory arena of size 64KB. Each survivor is loaded to a random address with a stack of 2,048 bytes and a full set of registers. The distance between two survivors and the arena’s edges is at least 1,024 bytes. The last survivor alive is the winner and gets one point. If several survivors stay alive, the point is divided equally between all. A survivor is disqualified if it runs an illegal command or attempts to access a memory address outside the arena or its stack. Each battle of the game runs for 200,000 rounds or until only one survivor is left, whichever comes first. In every round, the next command of each survivor is executed in a round-robin fashion. The order of the survivor’s execution is changed randomly for each game. The memory arena image and scoreboard that monitors the game’s progress are depicted in Figure 1. In most cases, each participant has two programs that collaborate together. The game includes the following special commands: WAITx4 increases the survivor’s speed, allowing it to run several opcodes in a single round, INT 0x86 implements “heavy bombing” by writing 256 bytes into memory, and INT 0x87 implements “smart bombing” by re-writing a pattern of 4 bytes.
@@
CodeGuru competition has taken place every year since 2005 among outstanding high-school students. Each year, the level rises, with more sophisticated survivors written. This work aims to evolve survivors that will win the top survivors of previous years by finding their weaknesses. Our goal is not to win the competition but rather to show that GP can be utilized for evolving code in an adversarial environment. Thus, we evolve a different survivor for each past survivor rather than evolving one survivor that takes them all.
4. Method
@@
4.1. Representation
Each individual consists of two programs, called parts (see Section 3), represented by an AST. The AST follows a grammar defined by a BNF—a meta-syntax notation for context-free grammar consisting of derivation rules. Since assembly is a symbolic programming language, it can be represented using it. The terminals are opcodes and operands, and the functions are structures in the language (see listings LABEL:tab:1_terminal_set and LABEL:tab:t2_function_set in the appendix). We defined our types and derivation rules based on assembly language constraints. For example, we defined an unary command as a command consisting of an opcode that takes only one operand. Notably, except for a few CodeGuru special operators, our BNF is general and can match any 8086 assembly code. There are assembly commands that are not supported by the game’s engine and were left out of the grammar to preserve legal programs.
4.2. Evaluation
We evaluated survivors by running a CodeGuru game of 200 battles against the selected human-written survivor. The game’s engine is an open-source Java program that outputs the final scores for each game. As previously explained, the score is one point given to the last survivor alive. If several survivors stay alive, the point is divided equally between all. We modified the engine to produce more information about each game, as elaborated by the fitness function that has four parts:
@@ 4.2. Evaluation
The first two parts encourage evolution to win the competitions and survive for longer periods (respectively). The last two parts encourage the evolution of programs that write in different memory places, which enhances the chance of damaging opponents. We refer to the score parameter as the most significant since it reflects the performance compared to the adversary. Nevertheless, the other parts are important for guiding the evolution towards the different sub-goals and discriminating the individuals. We also defined a bloat weight parameter, which equals $10^{-5}$. It slightly lowers the fitness of large evolved trees in order to prevent trees from bloating and yet allows large but powerful trees to evolve.
@@
It produces fitness values in the range of $[0,2.5]$, which does not produce sharp deviations.
4.3. Genetic Operators

### v2

@@
Abstract
In this work, we evolve Assembly code for the CodeGuru competition. The goal is to create a survivor—an Assembly program that runs the longest in shared memory, by resisting attacks from adversary survivors and finding their weaknesses. For evolving top-notch solvers, we specify a Backus Normal Form (BNF) for the Assembly language and synthesize the code from scratch using Genetic Programming (GP). We evaluate the survivors by running CodeGuru games against human-written winning survivors. Our evolved programs found weaknesses in the programs they were trained against and utilized them. To push evolution further, we implemented memetic operators that utilize machine learning to explore the solution space effectively. This work has important applications for cyber-security as we utilize evolution to detect weaknesses in survivors. The Assembly BNF is domain-independent; thus, by modifying the fitness function, it can detect code weaknesses and help fix them. Finally, the CodeGuru competition offers a novel platform for analyzing GP and code evolution in adversarial environments. To support further research in this direction, we provide a thorough qualitative analysis of the evolved survivors and the weaknesses found.
Keywords
@@ Abstract
1 Introduction
@@
CodeGuru Xtreme codeguru_repo is a coding competition where short 8086 Assembly programs, called survivors, are loaded into a random address in a virtual computer memory arena. Their goal is to defeat all other survivors by staying the last program to run. An opponent is defeated when it runs an illegal command caused, e.g., by overwriting its memory. A screen-shot of the game is depicted in Figure 1. Each survivor gets a different color in the arena, representing the bytes it wrote to the shared memory. We elaborate on the game in Section 3.
@@
Figure 1: The CodeGuru Xtreme game. On the left are the survivors; on the center is the code of the selected survivor; and on the right is the arena, i.e., the memory status. Each survivor gets a different color in the arena, representing its written bytes.
@@
In this work, we evolve winning survivors from scratch, i.e., from randomly generated Assembly code, and without access to the source code of known survivors. For this task, we utilize Grammar-Guided Genetic Programming (G3P)—an evolutionary computation technique that incorporates GP principles, employs context-free grammar and operates directly with tree-based representations. G3P allows us to evolve Assembly programs following grammar-type constraints. The goal of the individuals, embodied by the fitness function, is to overtake adversaries and win the game. The evolved code is represented using an Abstract Syntax Tree (AST) based on matching Assembly Backus Normal Form (BNF) we defined. A BNF is a meta-syntax notation for context-free grammar consisting of derivation rules. Since our BNF is general and domain-agnostic, the approach applies to generating Assembly programs for other domains and processors.
@@
No previous work has been done on the CodeGuru Xtreme game, except for an undergraduate project Darwin8086 and some early work on the “Core War” game Andersen2001TheGE; corno2003exploiting, which served as the basis CodeGuru Xtreme (see Section 3). As elaborated in Section 2, some work has been done on the evolution of low-level languages (i.e., Assembly and Java bytecode), and some work has been done to improve existing Assembly code. Note that improving existing code is a simpler task than generating new code from scratch, as the former’s state space is much smaller Petke2018GeneticImprovementSoftware; Banzhaf2018SomeRemarksCode.
@@
CodeGuru Xtreme competition has been running since 2005, with all past survivors publicly available. Thus, winning is considerably tricky and requires, among other qualities, a good understanding of the 8086 Assembly language.
This work combines a fitness approximation approach based on machine learning (ML) to overcome the expensive fitness evaluation. We utilize this approximation for developing memetic operators—genetic operators that incorporate local search to enhance the exploration, leading to significant improvements in the overall evolutionary computation.
@@
Our work has implications for cyber-security. We utilize evolution to detect and exploit weaknesses in other survivors. Furthermore, understanding the Assembly language is a necessity for some viruses. By modifying the fitness function, our approach can be used for detecting weaknesses in code and help in fixing them, detecting suspicious adversarial code, or, on the contrary, can be intended to avoid security mechanisms by mutating the virus to avoid detection while keeping its functionality. In Section 6, we use the CodeGuru Xtreme game and GP for mimicking Assembly viruses and demonstrating how our method can be used for avoiding anti-virus detection.
@@
The CodeGuru Xtreme competition provides a unique opportunity to analyze GP and code evolution in adversarial environments. In order to encourage further research in this area, we conducted a comprehensive qualitative analysis of the evolved survivors and identified their strengths and weaknesses. This analysis sheds light on the effectiveness of the evolved code and provides valuable insights for future improvements and advancements in the field.
The contributions of this work are:
  • We develop a generic Assembly BNF.
  • We demonstrate, for the first time, how evolution can evolve Assembly code in an adversarial environment, completely from scratch.
  • We provide a qualitative analysis of the evolved survivors and the weaknesses found to facilitate further research in the field of cyber-security.
  • We demonstrate how CodeGuru and our approach can be utilized to mimic and explore virus behavior.
2 Related Work
@@
Several works have been done on low-level code evolution, some similar to Assembly, like Verilog (a hardware description language) and Java bytecode.
ulya2005automatic used Grammatical Evolution (GE) for evolving a simple program of a one-bit full adder. In spite of the strong bias they employed, the success achieved was only about 5.7%.  orlov2011flight proposed a method for genetic improvement and repair of existing Java programs or any software that can be compiled into Java bytecode. Although Java bytecode resembles Assembly, it has a simplified representation and does not have direct memory access. This contrasts with Assembly, which has a very strong correspondence between its instructions, the architecture’s machine code instructions, and memory. Orlov and Sipper seeded the initial population with copies of a single hand-crafted individual and improved it over generations while we evolve Assembly code from scratch. Rosin2019Stepping synthesized simple programs with loops from input/output examples. They targeted a simplified low-level language similar to Assembly, where each instruction consists of an opcode and a single operand.
@@
Some works focused on constrained Assembly evolution—specific routines, predefined input and output tables, and manually written code parts. Serruto2017Automatic applied multi-objective linear GP for the automatic generation of specific Assembly driver routines. The evolved programs did not contain jump instructions because they could form infinite loops, in contrast to our wish to include them in the generated code. The results showed that the automatically generated microcontroller code for specific tasks can compete with a human programmer with a smaller code size or faster execution. We aim to recreate this result within an adversarial environment. Similarly, Ferrel2020Genetic presented a methodology for writing Arduino programs using an automatic generator of Assembly language routines based on a cooperative co-evolutionary multi-objective linear GP. They decomposed the problem into sub-components, generating about 73% of the program, and the remaining 27%, which are the main program and initial configuration routines, are manually written. In our case, we cannot break the goal of winning an opponent to sub-tasks without reading its code first, which we avoid. In addition, there is no clear way to represent a winning result using an input-output table. We need to evolve from scratch, a winning program, including jumps and loops, completely automatically.
@@
domonik2021recent surveyed recent developments in program synthesis with evolutionary algorithms and found that the most influential approaches in the field are stack-based GP (usually PushGP), G3P, and Linear GP. PushGP produces code only in the Push language and cannot be used outside of a Push interpreter, while G3P and Linear GP can produce source code in any language, including Assembly.
Pantridge2017On compared the synthesis capabilities of PushGP and G3P over different programming languages. It was done on problems of two types: the first is usually approached using machine code or Assembly language (basic execution models problems), while the second is usually approached using high-level languages (general program synthesis benchmark suite). In the low-level field, the best-synthesized program was in TerpreT—a probabilistic programming language designed for inductive program synthesis. It was able to solve all the basic execution model problems, in contrast to PushGP, which succeeded in 6 out of 8. Sadly, there were no gathered results for G3P in this field. In the high-level field, G3P was able to solve almost as many problems as PushGP in the software synthesis benchmark suite.
@@
GP usage in cyber-Security
Oreilly2020Adversarial stated that combining GP and competitive co-evolutionary algorithms enables evolving complex behaviors to be used as abstractions by adversaries. The paper presents the RIVALS framework, which serves as a testbed for computational modeling and simulation of the dynamics of networks under attack. Hu2020Optimal describe the attack-defense relationship using LQRD (Logit Quantal Response Dynamics). They describe how, by analyzing the evolutionary equilibrium, we can obtain the optimal defense strategy and demonstrate it on WannaCry malware.
In addition to the works that use GP to obtain strategies, others utilize it to evolve and modify malware. Noreen2009Evolvable developed a feature representation of the Bagel malware family and, using GA and a few training samples, were able to evolve new unknown variants. Castro2019AIMED used GP to evade malware detection by automatically finding code optimizations that, when injected into previously detected malware, result in misclassification of the malware scanner. They implemented their framework as a sandbox to track how the code samples behave. Murali2022Adapting used novelty search in order to generate malware variants of greater diversity to evade detection. They used Assembly code samples and represented them in linear or graph representations. Generic Assembly code transformation functions are applied as operators: inserting fake instructions forced jmp, unreachable blocks, and conditional jmps. They were able to evade over 98% of popular scanners using this technique.
In our work, we demonstrate the ability to evade detection using GP with general Assembly grammar and domain-independent operators. Unlike the described above works, we do not create special operators for code obfuscation. We utilize CodeGuru as an adversarial game framework to evolve Assembly code, which is able to evade an adversary and is specially designed to overtake it based on a characterizing signature. This resembles diverse forms of viruses that avoid malware scanners.
3 CodeGuru Xtreme
@@
CodeGuru Xtreme by codeguru_repo is a coding competition based on “Core War”—a 1984 programming game created by D. G. Jones and A. K. Dewdney dewdney1984recreational. In CodeGuru Xtreme, short 8086 Assembly programs (at most 512 bytes long) of 16-bit commands, called survivors, are loaded into random space on a virtual computer memory arena of size 64KB. Each survivor is loaded to a random address with a stack of 2,048 bytes and a full set of registers. The distance between two survivors and the arena’s edges is at least 1,024 bytes. The last survivor alive is the winner and gets one point. If several survivors stay alive, the point is divided equally between all. A survivor is disqualified if it runs an illegal command or attempts to access a memory address outside the arena or its stack. A survivor can be forced to run an illegal command due to a writing operation an opponent previously performed on its code or on an address it reads from. Therefore, performing various writings to memory enhances the chance of damaging opponents. Each battle of the game runs for 200,000 rounds or until only one survivor is left, whichever comes first. In every round, the next command of each survivor is executed in a round-robin fashion. The order of the survivor’s execution is changed randomly for each game. The memory arena image and scoreboard that monitors the game’s progress are depicted in Figure 1. In most cases, each participant has two programs, called parts, that can collaborate together. Each part has its own registers and is loaded into a different place in memory, although they both share a stack. The parts are executed separately, one after the other. Their scores are joined together at the end of the battle and create the survivor’s score. This allows the design of a survivor with two parts collaborating together via a shared stack or with two parts completely independently running; both designs are used as a force multiplier for maximizing the survivor’s abilities. The game includes the following special commands: WAITx4 increases the survivor’s speed, allowing it to run several opcodes in a single round, INT 0x86 writes 256 bytes into memory, and INT 0x87 re-writes a pattern of 4 bytes. The special commands allow performing several rounds of actions in one round.
@@
The CodeGuru competition has taken place every year since 2005 among outstanding high-school students. Each year, the level rises, with more sophisticated survivors written. This work aims to evolve survivors that will win the top survivors of previous years by finding their weaknesses. Our goal is not to win the competition but rather to show that GP can be utilized for evolving code in an adversarial environment. Thus, we evolve a different survivor for each past survivor rather than evolving one survivor that takes them all.
4 Method
@@
4.1 Representation
Each individual consists of two programs, called parts (see Section 3), represented by an AST that follows a grammar defined by a BNF. Since Assembly is a symbolic programming language, it can be represented using it. The terminals are opcodes and operands, and the functions are structures in the language (see listings LABEL:tab:1_terminal_set and LABEL:tab:t2_function_set in the appendix). We define our types and derivation rules based on Assembly language constraints. For example, we define an unary command as a command consisting of an opcode that takes only one operand. Notably, except for a few CodeGuru special operators, our BNF is general and can match any 8086 Assembly code. There are Assembly commands that are not supported by the game’s engine and were left out of the grammar to preserve legal programs.
4.2 Fitness Function
We evaluate survivors’ fitness by running a CodeGuru game of 200 battles with the selected human-written survivor the evolution performed against. The game’s engine is an open-source Java program that outputs the final scores for each game. As previously explained, the score is one point given to the last survivor alive. If several survivors stay alive, the point is divided equally between all. We modified the engine to produce more information about each game, as elaborated by the fitness function that has four parts:
@@ 4.2 Fitness Function
The first two parts encourage evolution to win the competitions and survive for longer periods (respectively). The last two parts encourage the evolution of programs that write in different memory places, which enhances the chance of damaging opponents. We refer to the score parameter as the most significant since it reflects the performance compared to the adversary. Nevertheless, the other parts are important for guiding the evolution towards the different sub-goals and discriminating the individuals. Division by 10 and $\log_{10}$ were used on the original values of lifetime and written_bytes in order to normalize them to an easy-to-process range yet maintain the tendency they represent. We also defined a bloat weight parameter, which equals $10^{-5}$. It slightly lowers the fitness of large evolved trees in order to prevent them from bloating and yet allows large but powerful trees to evolve.
@@
It produces fitness values in the range of $[0,2.5]$, which does not produce sharp deviations. The chosen weights reflect the above-elaborated goals. The score is doubled due to its significance, resulting in a value in the range of $[0,2]$. The additional parameters were multiplied by weights, resulting in a mutual sum of $0.5$ at most in order not to overshadow the score. According to conducted experiments, the writing_rate parameter resulted in the smallest survivor’s improvement, thus receiving the lowest weight. Experiments also showed that evolution was able to independently learn the importance of the survivor’s lifetime, on the contrary to the importance of performed writings, which it learned seldom. Thus, high and medium weights were given to written_bytes and lifetime parameters, respectively, to accelerate the improvement process.
4.3 Genetic Operators